he number of bacteria in a certain population increases according to a continuous exponential growth model, with a growth rate parameter of
3.9%
per hour. How many hours does it take for the size of the sample to double?
Note: This is a continuous exponential growth model.
Do not round any intermediate computations, and round your answer to the nearest hundredth.

Answer by stanbon(75887)

he number of bacteria in a certain population increases according to a continuous exponential growth model, with a growth rate parameter of 3.9% per hour.
How many hours does it take for the size of the sample to double?
Note: This is a continuous exponential growth model.
Do not round any intermediate computations, and round your answer to the nearest hundredth.
-----
A(t) = Ao*(1.039)^h
---------
Solve:: 1.039^h = 2
h = log(2)/(log(1.039)
h = 18.12 hours
-------------------------
Cheers,
Stan H.
---------
-----